Gaseous butane CH3CH22CH3 will react with gaseous oxygen O2 to produce gaseous carbon dioxide CO2 and gaseous water H2O

. Suppose 11.6 g of butane is mixed with 10. g of oxygen. Calculate the maximum mass of water that could be produced by the chemical reaction. Be sure your answer has the correct number of significant digits.


1
Expert's answer
2020-10-26T14:49:08-0400

0.3125mol            Xmol

2C4H10 + 13O2 = 8CO2 + 10H2O

            13mol               10mol

n(C4H10) = m(C4H10)/M(C4H10) = 11.6g/58g/mol = 0.2 mol (excess).

n(O2) = m(O2)/M(O2) = 10g/32g/mol = 0.3125mol (deficiency).

n(H2O) = X = (0.3125 × 10 / 13) mol = 0.2403mol.

m(H2O) = n(H2O) × M(H2O) = 0.2403 mol × 18 g/mol = 4.3254g ≈ 4.33g.